# CogDevelop2K: Reversed Cognitive Development in Multimodal Large Language Models

### v1

@@ 5 Conclusion
In this paper, we explored the cognitive capabilities of Multi-modal Large Language Models (MLLMs) through the lens of core cognitive abilities that underpin human intelligence. By introducing CogDevelop2K, a novel benchmark that spans 12 subconcepts across developmental stages, we aimed to assess the fundamental understanding and reasoning capacities of MLLMs. Our evaluation of 46 models revealed intriguing insights, including a reversed cognitive developmental trajectory compared to humans. This finding raises questions about whether MLLMs truly comprehend tasks or simply exhibit performance without genuine understanding. These results underscore the need for further investigation into the cognitive foundations of MLLMs, as well as the influence of evaluation strategies and prompting techniques in shaping their outcomes. Ultimately, this study serves as a step toward unraveling the nature of MLLM intelligence and their potential limitations in mirroring human cognitive development.

### v2

@@ 5 Conclusion
In this paper, we explored the cognitive capabilities of Multi-modal Large Language Models (MLLMs) through the lens of core cognitive abilities that underpin human intelligence. By introducing CogDevelop2K, a novel benchmark that spans 12 subconcepts across developmental stages, we aimed to assess the fundamental understanding and reasoning capacities of MLLMs. Our evaluation of 46 models revealed intriguing insights, including a reversed cognitive developmental trajectory compared to humans. This finding raises questions about whether MLLMs truly comprehend tasks or simply exhibit performance without genuine understanding. These results underscore the need for further investigation into the cognitive foundations of MLLMs, as well as the influence of evaluation strategies and prompting techniques in shaping their outcomes. Ultimately, this study serves as a step toward unraveling the nature of MLLM intelligence and their potential limitations in mirroring human cognition.